If the length of a square is increased by 100%, how the area of the square is increased by 400%? (please explain the meyhod)

Answer by Alan3354(69443)

If the length of a square is increased by 100%, how the area of the square is increased by 400%? (please explain the meyhod)
--------------
It isn't, so there's no explanation.
------------
eg, a square with sides of length 1.
Area = 1*1 = 1 sq unit
-------
Increasing the sides 100% --> sides = 2
Area = 2*2 = 4 sq units.
----
The increase is 4 - 1 = 3 sq units
Increase/original = 3/1 = 3
3 = 300% increase, not 400% increase.
======================
This display of ignorance surfaces freqently.
eg, 4 times more is not the same as 4 times as much.
4 times more is 5 times as much.
==============
I check the site apod.nasa.gov for the daily photos.
I wrote to the site administrators about saying that something moving "4 times faster" was the same as 4 times as much, but they didn't seem to get it.
----------------
Don't even think about saying "10 times smaller." 1 time smaller is zero.